OpenStudy (anonymous):

Can someone help me understand quadratic formula x^2+6x+4=0 is the example problem and I have no idea how they got the answer

OpenStudy (zarkon):

\[ax^2+bx+c=0 \Rightarrow x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] here your a=1, b=6 and c=4

OpenStudy (anonymous):



OpenStudy (zarkon):

\[x=\frac{-6\pm\sqrt{6^2-4\cdot1\cdot4}}{2\cdot 1}\]

OpenStudy (anonymous):

yea I just worte that but where does the one come from?

OpenStudy (zarkon):

\[x^2+6x+4=1\cdot x^2+6\cdot x+4\]

OpenStudy (zarkon):

it is next to the \(x^2\)

OpenStudy (anonymous):

ok so how do you dcide what goes on top and what goes on bottom